# Multi-person eye tracking for real-world scene perception in social settings

### v1

@@ 3.2.2. CentralCam
This module is responsible for handling a centrally mounted camera that records a stationary perspective of the shared scene viewed by all the eye tracking device wearers. The module supports off-the-shelf webcams connected via USB interface as well as network RTSP streams from cameras accessible wirelessly or via ethernet. The module grabs and timestamps independent frames from the stream. The module also computes dynamic frame-rate in frames per second (FPS), temporal jitter calculated as the mean and standard deviation of inter-frame time difference, and dropped frame counts. The final outputs from this module include the recorded video in an mp4 container, along with timestamps for each recorded frame, and the calculated metrics as comma-separated values (csv). Additionally, the module offers the option to save the RAW video as output. The RAW video format occupies much larger disk space, however, it requires much less computation (no decoding) and stores lossless data for post-hoc processing or correction. In the streaming mode, the module initialises a Kafka producer and pushes each incoming frame to the producer. Before sending it to the producer, the module resizes the frame, applies JPEG compression, and bundles it into a JSON object along with the recorded timestamp.
@@ 3.2.3. GlassesStream
This module implements the streaming of gaze and worldview data streams in the streaming mode of operation. The module initialises a Kafka producer, parses incoming gaze and video data streams from selected eye tracking devices on the network and publishes them to the Kafka cluster. The module also starts a metrics server, incorporating Grafana and Prometheus services, to monitor the Kafka cluster.
@@ 3.2.4. Flask Monitor
The Flask Monitor module provides real-time monitoring of the incoming data streams to verify device connectivity and data quality. The module initialises a Kafka consumer that reads each topic and updates corresponding graphs for that data stream. The module hosts a flask server that provides a graphical web-browser interface. Worldview streams from all devices are displayed on this interface with a configurable refresh rate (default 1Hz) and incoming frame rate (in FPS) which is used to colour code the display window of each device. For example, if the current sample rate of an incoming worldview stream drops to zero, its display window boundary turns red. Incoming gaze is displayed as a dynamically updating time-series of x and y coordinates, with a configurable refresh rate. The sample rates for gaze and worldview streams are set to a low value by default to reduce stress on the server and keep the monitor interface responsive. The incoming centralview is also consumed in a separate process and displayed in a separate window, similar to the worldview frames.
@@ 3.2.5. Homography
Since the gaze and worldviews from each pair of eye tracking glasses are egocentric to the wearer’s perspective, a direct comparison of spatial gaze measures from different wearers is not possible. The homography module maps gaze from egocentric worldview coordinate space to centralview coordinate space simultaneously for multiple devices. Incoming gaze samples (x, y coordinates) and worldview frames from each device are paired together with the nearest centralview frame using the marked timestamps from the respective input modules of each data stream (see 3.2). The module accommodates varying data sampling rates of the three data streams, since each stream’s sampling resolution depends upon factors like hardware configuration, recording parameters, and network bandwidth. The final temporal resolution of the homography module output is the same as the sampling rate of the provided centralview (i.e., for each centralview frame, the homography module attempts to map all gaze points from the provided worldviews+gaze pairs to that centralview frame). This mapping is accomplished by finding robust feature point correspondences between each worldview-centralview pair. Matched key points are then used to calculate a homography matrix using a least squares method and the computed matrix is finally used to transform gaze from worldview to centralview coordinate space. We apply a state-of-the-art deep learning model–SuperGlue (Sarlin et al., 2020)–to find robust feature point matches in the recording mode of operation in our Utility Test (see 4); however, the real-time operation of this module is dependent on the applied use-case, available computational resources, and future usability testing.
@@ 3.2.6. Visualisation
The visualisation module synchronises timestamped data from recorded worldview frames, gaze samples, and centralview frames with an additional input of the transformed gaze in the centralview space–computed by the homography module. This synchronised data can be used for visual analytics by presenting it through various visualisation techniques. We propose and implement the following visualisation examples in our Utility Study:
@@ 3.2.7. Analysis
The Analysis module provides methods to preprocess and evaluate the recorded egocentric gaze data as well as the homography-transformed gaze data for collective gaze analysis metrics. We propose and implement the following metrics that include traditional eye-tracking measures, such as heatmap similarity and correlation, scaled up to multi-person use cases, and novel collective measures, such as convex hull area, to quantify gaze dynamics in social group settings:
@@
  • Heatmap Similarity (SIM) and Correlation (CC): Similarity is calculated as the histogram intersection between the distributions of two heatmaps (ranges between 0 and 1 where 1 represents identical distributions) and correlation is represented by the Pearson correlation coefficient between them to quantify the degree of linear relationship between their pixel intensities.
@@
  • Convex Hull Area: Convex Hull Area provides a dynamic measure of gaze dispersion over time. It is calculated as the area of the smallest convex polygon that encloses all provided gaze samples at a given time point. The convex hull area was initially adapted to eye tracking analysis by Goldberg& Kotval (1999).
@@ 4. Utility Test
To test the applicability and scalability of our proposed framework with a large audience in a real-world context, we implemented the framework during a public event featuring a musical concert and a film screening. The event was held at the Large Interactive Virtual Environment (LIVELab) at McMaster University, Canada, which is a custom-built research concert hall with fully customisable spatial audio and room acoustics. The LIVELab holds 106 audience members; we had mobile eye-tracking glasses for 30 audience members. Since multi-person eye-tracking at such a scale, with a participant sample size of 60 (across two event days), has never been attempted before, we detail our choices for the system design components and evaluate our proposed analysis metrics to derive insights from multi-person eye movement data.
4.1. Methods
@@ 4.1.1. Apparatus
Eye tracking glasses: We employed Pupil Labs Neon eye tracking devices (Baumann and Dierkes, 2023) that apply a deep learning pipeline to provide reliable gaze predictions in unrestricted settings, requiring minimal calibration and setup. At the core of these eye trackers is a hardware module equipped with binocular infrared cameras that capture eye images with a resolution of 192 x 192 pixels at 200 Hz and a front-facing RGB camera that records the worldview with a resolution of 1600 x 1200 pixels at 30 Hz. The module is mounted to a 3D-printed, wearable frame that connects to an accompanying Android smartphone (Motorola Edge 40 Pro). The minimal form factor of the frames is designed to be unobtrusive and effortless for participants, making them a good fit for longer-duration studies. The accompanying Android phone runs a proprietary vendor application–“Neon Companion”–that hosts a suite of HTTP REST Application Programming Interfaces (APIs) enabling remote streaming of data and controlling operations such as start/stop recording. During the data collection event, each pair of eye tracking glasses was connected to the companion smartphone device running Pupil’s Labs NEON Companion android application in the background. The phones were locked and connected to an Anker 6-in-1 USB-C hub that extended the ports to connect them to ethernet, power, and the NEON glasses concurrently. Each smartphone was secured in a bag attached to the respective participant’s chair, ensuring free movement of the seated participant (see Figure 3B). The ethernet cable connected the device to a network switch placed under the false floor, in the centre of the audience rows. The global NTP server for all smartphones, the central camera, and the server computers were all set to the University’s NTP server available on the local network.
Figure 3. Study setup. A represents the stage setup during the concert part of the event. The film was presented on the video wall visible behind the performers. During the film, the stage was cleared and/or covered. B shows audience chairs with eye tracking glasses attached to each seat. The eye tracking glasses were connected to companion smartphones that were secured in pockets attached on either side of each pair of seats.
@@
Central view camera: A PTZOptics pt30x-sdi-g2 camera was mounted central to the stage, positioned behind the audience seating. The camera captured an unobstructed view of the stage with a resolution of 720x480 px at 60 Hz. The camera feed was streamed on the network via Real-Time Streaming Protocol with H264 encoding. Figure 3A demonstrates an example frame from the central camera.
@@
Network Infrastructure: Each pair of Neon glasses requires a bandwidth of 6-8 Megabytes per second (MBps) to stream gaze and worldview data. For real-time streaming operations, it is critical to consider the available bandwidth on a network to identify how many concurrent eye tracking glasses’ data can be streamed. In this test, we used the recording mode of operation which has lower bandwidth requirements, since real-time data streams do not need to be accessed on the server. Nonetheless, we leveraged the stationary positioning of participants in our setup to provide precise timing information and optimal network reliability by connecting each eye tracking device to a network switch (TP Link TL-SG1048) via ethernet. The network did not have internet access, to disable any background network activity on the connected Android smartphones.
@@
4.1.2. Implementation of proposed system design in Recording mode
GlassesRecord: We implemented the GlassesRecord module to be compatible with the Neon glasses using the pupil-labs-real-time-api python library (https://pupil-labs-realtime-api.readthedocs.io) to send start, stop, discard, and save triggers for a recording to each device. Additionally, since the accompanying smartphones use Android as operating system, we used Android Debug Bridge (ADB) to retrieve status metrics including battery level, storage level, network ping, ADB server status, and connected USB devices (see Figure 4A). The status metrics facilitated continuous monitoring of devices and recording status. ADB was also used to remotely control the Android smartphones for remote troubleshooting and verifying indicators of failed or corrupted recordings that were not reported by the vendor application. These operations were important to prevent data loss and ensure that we would never need to interrupt the live social event to access a participants’ phone in the event of an error. The module operations were interfaced via an intuitive Terminal User Interface (TUI) built using the python Textual library (https://textual.textualize.io/), providing low runtime requirements and cross platform operation. The TUI (see Figure 4) displays a list of all eye tracking devices found on the local network, along with accompanying colour-coded status fields for each device. An action panel at the bottom displays a list of operations (actions) that can be performed on selected devices by pressing the corresponding action key. Multiple devices can be selected at a time on the TUI, including hotkey actions of Spacebar to select a single device; Shift + Arrow Up/Down: to expand selection up/down; Ctrl+A to select all devices; and Exclamation mark (!) to reverse selection.
Figure 4. Data recording and processing interfaces A. Terminal User Interface (TUI) for the GlassesRecord module used to monitor and operate all 30 devices concurrently. B Command Line Interface for offline analysis and visualisation of the recorded data.
@@ 4.1.2. Implementation of proposed system design in Recording mode
Homography, Analysis, and Visualisation: We applied a state-of-the-art deep learning model– SuperGlue (Sarlin et al., 2020)–to find robust feature point correspondences in our Homography module. The analysis, and visualisation modules’ methods were implemented in Python 3.12. Blinks were identified using the Neon blink-detection algorithm and filtered out of the raw gaze data before analysis and visualisation. As with the GlassesRecord module, we developed an interactive command line utility using the Python questionary (questionary.readthedocs.io) library (see Figure 4B) to interact with the three modules. The offline utility offers formatting of the directory structure for compatibility with the modules, selecting a subset of recordings to perform operations, such as homography, visualisation, gaze metrics calculation, etc., and storing output to the local filesystem.
@@ 4.1.3. Stimuli
Music Performance: American percussionist-composers, Allen Otte and John Lane, performed the Canadian premiere of The Innocents – an hour-long performance art piece centred on the issue of wrongful incarceration. The work might best be described as social justice advocacy through performance art. The hour-long composition consists of home-made instruments, spoken texts, and electronic soundscapes. Across 17 tableaux–many of which elide into one another–the audience is presented with dramatic soundscapes that explore issues related to the American criminal justice system, such as mistaken identity, injustice, truth, politics, psychology, and resilience. Some tableaux are too long, uncomfortable, confusing; others are direct, melodic, and in familiar styles.
Film: Otte and Lane’s The Innocents is now the subject of an award-winning feature-length (1hr 20min) documentary film, directed by Wojciech Lorenc. The film showcases Otte and Lane’s compositional process and friendship, following them on the road as they tour the United States performing. On tour, we witness conversations with relevant stakeholders, such as lawyers, non-profit organisations, students, and exonerees. Woven throughout the film is the story of exoneree Anna Vasquez who served 13 years in prison for a crime that never occurred. We witness Anna tell parts of her story and ultimately her reaction to watching Otte and Lane’s live performance.
@@ 4.1.4. Study Design
During the event, in addition to wearing eye-tracking glasses, participants also wore a smart watch to monitor cardiac activity and completed a series of surveys. Parallel to this in-person event, we also presented a livestream online. We recorded all of the same measures (eye, heart, survey), via web browser and webcam ((Saxena et al., 2022, 2023). We were interested in the effect social co-presence might have on our measures of interest. Detailed analyses related to our cognitive neuroscientific and social psychological questions will be reported elsewhere. In the current study, we focus on validating the utility of our proposed system design.
@@ 4.1.5. Procedure
Before the event, ADB connection was established between the server and each smartphone device. A force synchronisation to the assigned NTP server on the smartphones was also done. Participants were seated evenly across rows two to four in the audience, hereby referred to as participant-rows one to three in this paper. After taking their seat, participants were instructed on how to put the glasses on. Trained research assistants completed a quick calibration procedure to correct any static gaze offsets on the Neon Companion application. Participants retained their seating positions over the entire study duration.
To ensure proper connection to all eye tracking devices, test recordings for a short length were made before the actual recording session was started. We started the recording session before the event host came on stage. After the start of the event, no interruptions were allowed and all interactions with the recording devices were done remotely using the GlassesRecord module TUI to trigger recordings and monitor the eye tracking devices (see Figure 4A). The CentralCam module was used to trigger recording on/off on the central camera.
In between film and performance (or vice versa), there was an intermission of 30 minutes, during which we stopped recording and participants took off their glasses, completed a short survey, went to the lobby, used the restroom, etc. The central camera recording was not stopped during event intermission. After intermission, the same calibration procedure was repeated before the second half of the event started. An additional calibration procedure where participants followed a moving calibration target on the videowall with their eyes was also conducted before the second half. Recordings were stopped when the host came onstage at the end of the second half to close the event. The event started at 7 pm and was, on average, 3 hours and 15 minutes long for each of the two days (including intermission).

### v2

@@ 3.2.3. CentralCam
The CentralCam module manages a centrally mounted camera that captures a stationary view of the shared scene. This camera can be interfaced as a USB peripheral or a network device via Real-Time Streaming Protocol (RTSP). The module grabs and timestamps independent frames from the incoming camera stream. It also computes peformance metrics such as dynamic frame-rate in frames per second (FPS), temporal jitter calculated as the mean and standard deviation of inter-frame time difference, and dropped frame counts. In recording mode, the module offers flexibility to store the recorded video in an mp4 container or a raw format. The raw format, while consuming more storage space, requires less computational overhead since it avoids decoding and preserves lossless data for later processing or correction. In the streaming mode, the module initialises a Kafka producer and pushes each incoming frame and its timestamp to the producer.
@@ 3.2.4. GlassesStream
GlassStream implements the streaming of gaze and worldview data streams in the streaming mode of operation. The module initialises a Kafka producer, parses incoming gaze and video data streams from selected eye-tracking devices on the network and publishes them to the Kafka cluster. The module also initiates a metrics server, incorporating Grafana and Prometheus services, to monitor the Kafka cluster.
@@ 3.2.5. Flask Monitor
The Flask Monitor module provides a real-time display of the incoming data streams to verify data integrity. The module initialises a Kafka consumer that reads each topic and updates corresponding graphs for that data stream. The module hosts a flask web-application that displays worldview streams from all devices with a configurable refresh rate (default 1Hz) and incoming frame rate (in FPS). Incoming gaze is displayed as a dynamically updating time-series of x and y coordinates, with a configurable refresh rate. The sample rates for gaze and worldview streams are set to a low value by default to reduce stress on the server and keep the monitor interface responsive. The incoming centralview is consumed in a separate process and displayed in a separate window, similar to the worldview frames. The module interface also features alert mechanisms to notify when data flow ceases or the frame rate drops.
@@ 3.2.6. Homography
The homography module implements semantic gaze mapping (as defined in our approach 1) from egocentric worldview coordinate space to centralview coordinate space simultaneously for multiple devices. Applying a pin-hole camera model, the worldview captured from each wearer’s perspective can be assumed to lie on a planar surface at an infinite distance. The projective transformation between the worldview and centralview for the same scene is then mathematically related by a planar homography. Incoming gaze samples (x, y coordinates) and worldview frames from each device are paired together with the nearest centralview frame using the marked timestamps from the respective input modules of each data stream. The module accommodates varying data sampling rates of the three data streams, since each stream’s sampling resolution depends upon factors like hardware configuration, recording parameters, and network bandwidth. The final temporal resolution of the homography module output is the same as the sampling rate of the provided centralview (i.e., for each centralview frame, the homography module attempts to map all gaze points from the provided worldviews+gaze pairs to that centralview frame). This mapping is accomplished by finding robust feature point correspondences between each worldview-centralview image pair. Matched key points are then used to calculate a homography matrix that is finally used to transform gaze from worldview to centralview coordinate space.
The performance of this module depends significantly on detecting and tracking robust feature point correspondences in dynamic, real-world social settings which often involve challenging artefacts caused by illumination changes, motion distortion, low-textured regions, and occlusions. In recording mode, the module can utilize large deep learning models, such as the state-of-the-art SuperGlue (Sarlin et al., 2020) model demonstrated in our Utility Test (see 4), thanks to the absence of computational and time constraints. However, for real-time operation in streaming mode, simpler algorithms like Bay et al. (2008); Rublee et al. (2011); Tomasi and Kanade (1991) are preferred, balancing the trade-off between speed and accuracy.
@@ 3.2.7. Visualisation
The visualisation module aligns timestamped worldview frames, gaze samples, and centralview frames with an additional input of the transformed gaze (in the centralview space) computed by the homography module. We design the following visualisation modes in this module:
@@ 3.2.8. Analysis
The Analysis module provides methods to preprocess and evaluate the recorded egocentric gaze data as well as the homography-transformed gaze data for collective gaze analysis metrics. We design the followign metrics including traditional eye-tracking measures, such as heatmap similarity and correlation, scaled up to multi-person use cases, and novel collective measures, such as convex hull area, to quantify gaze dynamics in social group settings:
@@
  • Heatmap Similarity (SIM) and Correlation (CC): SIM is calculated as the histogram intersection between the distributions of two heatmaps (ranges between 0 and 1 where 1 represents identical distributions) and CC is represented by the Pearson correlation coefficient between them to quantify the degree of linear relationship between their pixel intensities.
@@
  • Convex Hull Area: Convex Hull Area provides a dynamic measure of gaze dispersion over time. It is calculated as the area of the smallest convex polygon that encloses all provided gaze samples at a given time point. The convex hull area was initially adapted to eye-tracking analysis by Goldberg& Kotval (1999).
@@ 4. Utility Test
To test the viability and scalability of our framework with a large audience in a real-world context, we implemented the framework during a public event featuring a musical concert and a film screening. The event was held at the Large Interactive Virtual Environment (LIVELab) at McMaster University, Canada, which is a custom-built research concert hall with fully customisable spatial audio and room acoustics. The venue holds 106 audience members; we had mobile eye-tracking glasses for 30 audience members. Below, we detail our component choices for the implementation of SocialEyes in recording mode and provide results from the homography, analysis and visualisation modules to derive insights from multi-person eye movement data.
4.1. Method
@@ 4.1.1. Apparatus
  • Eye-tracking equipment: We employed Pupil Labs Neon eye-tracking devices (Baumann and Dierkes, 2023) that apply a deep learning pipeline to provide reliable gaze predictions in unrestricted settings, requiring minimal calibration and setup. At the core of these eye trackers is a hardware module equipped with binocular infrared cameras that capture eye images with a resolution of 192 x 192 pixels at 200 Hz and a front-facing RGB camera that records the worldview with a resolution of 1600 x 1200 pixels at 30 Hz. The module is mounted to a 3D-printed, wearable frame that connects to an accompanying Android smartphone (Motorola Edge 40 Pro). The minimal form factor of the frames is designed to be unobtrusive and effortless for participants, making them a good fit for longer-duration studies. The accompanying Android phone runs a proprietary vendor application–“Neon Companion”–that hosts a suite of HTTP REST Application Programming Interfaces (APIs) enabling remote streaming of data and controlling operations such as start/stop recording. During the data collection event, each pair of eye-tracking glasses was connected to the companion smartphone device running Pupil’s Labs NEON Companion android application in the background. The phones were locked and connected to an Anker 6-in-1 USB-C hub that extended the ports to connect them to ethernet, power, and the NEON glasses concurrently. Each smartphone was secured in a bag attached to the respective participant’s chair, ensuring free movement of the seated participant (see Figure 3B). The ethernet cable connected the device to a network switch placed under the false floor, in the centre of the audience rows. The global NTP server for all smartphones, the central camera, and the server computers were all set to the University’s NTP server accessible on the local network.
    Figure 3. Study setup. A represents the stage setup during the concert part of the event. The film was presented on the video wall visible behind the performers. During the film, the stage was cleared and/or covered. B shows audience chairs with eye-tracking glasses attached to each seat. The eye-tracking glasses were connected to companion smartphones that were secured in pockets attached on either side of each pair of seats.
@@
  • Centralview camera: A PTZOptics pt30x-sdi-g2 camera was mounted central to the stage, positioned behind the audience seating. The camera captured an unobstructed view of the stage with a resolution of 720x480 px at 60 Hz. The camera feed was streamed on the network via RTSP with H264 encoding. Figure 3A demonstrates an example frame from the central camera.
@@
  • Network Infrastructure: Each pair of Neon glasses requires a bandwidth of 6-8 Megabytes per second (MBps) to stream gaze and worldview data. For real-time streaming operations, it is critical to consider the available bandwidth on a network to identify how many concurrent eye-tracking glasses’ data can be streamed. In this test, we used the recording mode of operation which has lower bandwidth requirements, since real-time data streams do not need to be accessed on the server. Nonetheless, we leveraged the stationary positioning of participants in our setup to provide precise timing information and optimal network reliability by connecting each eye-tracking device to a network switch (TP Link TL-SG1048) via ethernet. The network did not have internet access, to disable any background network activity on the connected Android smartphones.
@@
4.1.2. SocialEyes implementation in Recording mode
  • GlassesRecord: We implemented the GlassesRecord module utilising the pupil-labs-real-time-api python library (https://pupil-labs-realtime-api.readthedocs.io) to send start, stop, discard, and save recording-triggers to each Neon device. Periodic clock offset measurements were recorded every ten seconds. Additionally, since the accompanying smartphones use Android as operating system, we used Android Debug Bridge (ADB) to retrieve status metrics including battery level, storage level, network ping, ADB server status, and connected USB devices (see Figure 4A). The status metrics facilitated continuous monitoring of devices and recording status. ADB was also used to remotely control the Android smartphones for remote troubleshooting and verifying indicators of failed or corrupted recordings that were not reported by the vendor application. These operations were important to prevent data loss and ensure that we would never need to interrupt the live social event to access a participants’ phone in the event of an error. The operator interface for this module was implemented as a Terminal User Interface (TUI) built using the python Textual library (https://textual.textualize.io/), featuring low runtime requirements and cross platform operation. The TUI (see Figure 4) displyed a list of all eye-tracking devices on the local network, along with accompanying colour-coded status fields for each device. An action panel at the bottom displayed a list of operations (actions) that can be performed on selected devices by pressing the corresponding action key. The TUI allowed selecting multiple devices to perform an action, including hotkey actions of Spacebar to select a single device; Shift + Arrow Up/Down: to expand selection up/down; Ctrl+A to select all devices; and Exclamation mark (!) to reverse selection.
    Figure 4. Data recording and processing interfaces. A Terminal User Interface (TUI) for the GlassesRecord module used to monitor and operate all 30 devices concurrently. B Command Line Interface to operate offline analysis and visualisation of the recorded data.
@@ 4.1.2. SocialEyes implementation in Recording mode
  • Homography, Analysis, and Visualisation: We applied a pre-trained Graph Neural Network–-SuperPoint/ SuperGlue (Sarlin et al., 2020)–to find robust feature point correspondences in our Homography module. Blinks were identified using the Neon blink-detection algorithm and filtered out of the raw gaze data before homography, analysis and visualisation. The operator interface for these modules was implemented as a command-line utility using the Python questionary (questionary.readthedocs.io) library (see Figure 4B). The command-line utility offered formatting of the directory structure for compatibility with modules functions, selecting a subset of recordings to perform functions, such as homography, visualisation, gaze metrics calculation, etc., and storing output to the local filesystem.
@@ 4.1.3. Stimuli
  • Music Performance: American percussionist-composers, Allen Otte and John Lane performed the Canadian premiere of The Innocents – hour-long performance art piece centred on the issue of wrongful incarceration. The work might best be described as social justice advocacy through performance art. The hour-long composition consists of home-made instruments, spoken texts, and electronic soundscapes. Across 17 tableaux–many of which elide into one another–the audience is presented with dramatic soundscapes that explore issues related to the American criminal justice system, such as mistaken identity, injustice, truth, politics, psychology, and resilience. Some tableaux are too long, uncomfortable, confusing; others are direct, melodic, and in familiar styles.
  • Film: Otte and Lane’s The Innocents] is now the subject of an award-winning feature-length (1hr 20min) documentary film, directed by Wojciech Lorenc. The film showcases Otte and Lane’s compositional process and friendship, following them on the road as they tour the United States performing. On tour, we witness conversations with relevant stakeholders, such as lawyers, non-profit organisations, students, and exonerees. Woven throughout the film is the story of exoneree Anna Vasquez who served 13 years in prison for a crime that never occurred. We witness them tell parts of their story and ultimately their reaction to watching Otte and Lane’s live performance.
@@ 4.1.4. Study Design
During the event, in addition to wearing eye-tracking glasses, participants also wore a smart watch to monitor cardiac activity and completed a series of surveys. Parallel to this in-person event, we also presented a livestream online. We recorded eye measures, via web browser and webcam ((Saxena et al., 2022, 2023)). We were interested in the effect social co-presence might have on our measures of interest. Detailed analyses related to our cognitive neuroscientific and social psychological questions will be reported elsewhere. In the current study, we focus on validating the utility of our software framework–SocialEyes.
@@ 4.1.5. Procedure
To ensure proper connection to all eye-tracking devices, test recordings for a short length were made before the actual recording session was started. We started the recording session before the event host came on stage. After the start of the event, no interruptions were allowed and all interactions with the recording devices were done remotely using the GlassesRecord module TUI to trigger recordings and monitor the eye-tracking devices (see Figure 4A).
In between film and performance (or vice versa), there was an intermission of 30 minutes, during which we stopped recording and participants took off their glasses, completed a short survey, went to the lobby, used the restroom, etc. The central camera recording was not stopped during event intermission. After intermission, the same calibration procedure was repeated before the second half of the event started. Recordings were stopped when the host came onstage at the end of the second half to close the event. The event started at 7 pm and was, on average, 3 hours and 15 minutes long for each of the two days (including intermission).